OpenStudy (butterflyhope):

Urgent *WILL MEDAL*: James invested $5000 in a money market account that earns 10% interest compounded monthly. Which equation can be used to find the balance in James's account at the end of 4 years? Thanks :))

OpenStudy (anonymous):

ok, so the formula of this equation is A=P(1+R/N)^nt

OpenStudy (anonymous):

A is the ending amount

OpenStudy (anonymous):

P is the amount invested

OpenStudy (anonymous):

r is the rate

OpenStudy (anonymous):

n is the number of compoundings per year

OpenStudy (anonymous):

t is total number of years

OpenStudy (anonymous):

If its monthly n=4

OpenStudy (anonymous):

now we substitute A=5000(1+.1/4)^16

OpenStudy (anonymous):

1.25^16 is ~35.5 Then we multiply by 5000

OpenStudy (anonymous):

So around 177k. He made a really good investment lol ;)

OpenStudy (anonymous):

OH wait. I was wrong. equation is A=5000(1+ .010/4)^16

OpenStudy (anonymous):

that would be around 1.17 x 5000

OpenStudy (anonymous):

Final answer is 5863. Good luck :)